# Generalized Random Forests using Fixed-Point Trees

### v3

@@ B.3 Proofs for Section 4
  • For a fixed parent node $P$, let $x_{P}$ denote the center of mass of the $X_{i}\in P$, and $r\coloneqq\sup_{\{i:X_{i}\in P\}}\left\lVert X_{i}-x_{P}\right\rVert$. We consider a limit where $n_{C_{j}}\to\infty$ and $r\to 0$, corresponding to small regions of $\mathcal{X}$ that identify small neighbourhoods of $\theta^{*}(x)$.
  • Let $\tilde{\theta}^{*}_{C_{j}}(x_{P})$ denote the pseudo-forest quantity defined in Athey et al. (2019). This is an unknown quantity, involving the underlying $\theta^{*}(\cdot)$ and $\nu^{*}(\cdot)$ functions, defined as
    $\tilde{\theta}^{*}_{C_{j}}(x_{P})\coloneqq\theta^{*}(x_{P})-\frac{1}{n_{C_{j}}}\xi^{\top}V(x_{P})^{-1}\sum_{\{i:X_{i}\in C_{j}\}}\psi_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(O_{i}),$ (34)
    where $V(x)$ is as defined in Assumption 2. Equivalently,
    $\tilde{\theta}^{*}_{C_{j}}(x_{P}):=\theta^{*}(x_{P})-\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}^{*}(x_{P}),$
    where $\rho_{i}^{*}(\cdot)$ are the influence functions/population pseudo-outcomes defined in Regularity Condition 1.
@@
Under the conditions of Proposition 4.1, the one-step fixed-point approximation $\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}$ and the pseudo-forest quantity $\tilde{\theta}^{*}_{C_{j}}(x_{P})$ defined in (34) are coupled at the following rate:
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}=\tilde{\theta}^{*}_{C_{j}}(x_{P})+o_{P}\left(r,1/\sqrt{n_{C_{j}}}\right).$
Proof.
Consider the difference $\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}-\tilde{\theta}^{*}_{C_{j}}(x_{P})$ and expand each term according to their definitions,
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}-\tilde{\theta}^{*}_{C_{j}}(x_{P})$ $=\hat{\theta}_{P}-\frac{1}{n_{C_{j}}}\xi^{\top}\sum_{\{i:X_{i}\in C_{j}\}}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$
$\hphantom{.....}-\theta^{*}(x_{P})+\frac{1}{n_{C_{j}}}\xi^{\top}V(x_{P})^{-1}\sum_{\{i:X_{i}\in C_{j}\}}\psi_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(O_{i}),$
[…]
$-\frac{1}{n_{C_{j}}}\xi^{\top}V(x_{P})^{-1}\sum_{\{i:X_{i}\in C_{j}\}}\left(\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})-\psi_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(O_{i})\right)=o_{P}(r).$
[…]
$-\frac{1}{n_{C_{j}}}\xi^{\top}\left(\mathbb{I}-V(x_{P})^{-1}\right)\sum_{i\in C_{j}}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$ $=\frac{1}{n_{C_{j}}}O_{P}(1)O_{P}\left(r,1/\sqrt{n_{C_{j}}}\right),$
$=o_{P}\left(r,1/\sqrt{n_{C_{j}}}\right),$
where $-\frac{1}{n_{C_{j}}}\xi^{\top}\left\lVert\mathbb{I}-V(x_{P})^{-1}\right\rVert=o_{P}(1)$, and where the bound
$\sum_{\{i:X_{i}\in C_{j}\}}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})=O_{P}\left(r,1/\sqrt{n_{C_{j}}}\right)$
is a bound established in Lemma 4 of Athey et al. (2019). Therefore,
$\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}-\tilde{\theta}^{*}_{C_{j}}(x_{P})=o_{P}\left(r,1/\sqrt{n_{C_{j}}}\right),$
@@
Under the conditions of Proposition 4.1, the one-step fixed-point approximation $\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}$ and the true local solution $\hat{\theta}_{C_{j}}$ are coupled at the following rate:
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}=\hat{\theta}_{C_{j}}+o_{P}\left(r,1/\sqrt{n_{C_{j}}}\right).$
Proof.
From Lemma B.2, we know that
$\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}=\tilde{\theta}^{*}_{C_{j}}(x_{P})+o_{P}\left(r,1/\sqrt{n_{C_{j}}}\right),$
@@
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}-\hat{\theta}_{C_{j}}=\tilde{\theta}^{*}_{C_{j}}(x_{P})-\hat{\theta}_{C_{j}}+o_{P}\left(r,1/\sqrt{n_{C_{j}}}\right).$
The expression $\tilde{\theta}^{*}_{C_{j}}(x_{P})-\hat{\theta}_{C_{j}}$ does not involve any fixed-point quantities and retains the rate argued in Proposition 2 of Athey et al. (2019):
$\tilde{\theta}^{*}_{C_{j}}(x_{P})-\hat{\theta}_{C_{j}}=o_{P}(r,1/\sqrt{n_{C_{j}}}).$
Therefore,
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}-\hat{\theta}_{C_{j}}=o_{P}\left(r,1/\sqrt{n_{C_{j}}}\right),$
as desired. ∎
Lemma B.4.
Under the conditions of Proposition 4.1,
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{1}}-\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{2}}=O_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right).$
Proof.
From Lemma B.3, we know that
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}-\hat{\theta}_{C_{j}}=o_{P}\left(r,1/\sqrt{n_{C_{j}}}\right),$
and hence
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{1}}-\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{2}}=\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}+o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right).$
Each $\hat{\theta}_{C_{j}}$ is $O_{P}(r)$-consistent for $\theta^{*}(x_{P})$ by Lemma 4 and Proposition 2 of Athey et al. (2019), and so
$\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}=O_{P}(r).$
Therefore
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{1}}-\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{2}}$ $=O_{P}(r)+o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right),$

### v4

@@ B.2 Proofs for Section 4
  • For a fixed parent node $P$, denote by $x_{P}$ the center of mass of the $X_{i}\in P$, and let $r\coloneqq\sup_{\{i:X_{i}\in P\}}\left\lVert X_{i}-x_{P}\right\rVert$ denote the radius of the parent $P$. Throughout, we consider an asymptotic regime where $n_{C_{j}}\to\infty$ and $r\to 0$, corresponding to leaves over $\mathcal{X}$ of vanishing radius. Further, $r$ and $n_{C_{j}}$ are related under the conditions of GRF Proposition 1, namely, $r^{-2}\ll n_{C_{j}}$ and hence $n_{C_{j}}r^{2}\to\infty$ and $1/\sqrt{n_{C_{j}}}=o(r)$.
  • Let $\theta_{C_{j}}^{*}$ denote the true parameter expectation over the child node:
    $\theta_{C_{j}}^{*}\coloneqq\mathbb{E}[\theta^{*}(X)\mid X\in C_{j}],\qquad j=1,2,$ (43)
    and let $\tilde{\theta}^{*}_{C_{j}}(x_{P})$ denote an oracle version of the gradient-based leaf statistic:
    $\tilde{\theta}^{*}_{C_{j}}(x_{P})\coloneqq\theta^{*}(x_{P})-\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\xi^{\top}V(x_{P})^{-1}\psi_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(O_{i}),$
    where $V(x)$ is the underlying local Jacobian in Assumption 2. Equivalently, in terms of the oracle pseudo-outcome/influence function $\rho_{i}^{*}(\cdot)$ defined in Regularity Condition 1,
    $\tilde{\theta}^{*}_{C_{j}}(x_{P})\coloneqq\theta^{*}(x_{P})+\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}^{*}(x_{P}).$
@@
Suppose Assumptions A.1 and Specifications A.2 hold. Then,
$\Delta(C_{1},C_{2})=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*}\right\rVert^{2}+o_{P}\left(r^{2},\frac{1}{n_{C_{1}}},~{}\frac{1}{n_{C_{2}}}\right).$
Proof of Lemma B.3.
Write the difference $\hat{\theta}_{C_{j}}-\theta_{C_{j}}^{*}$ as
$\hat{\theta}_{C_{j}}-\theta_{C_{j}}^{*}=\underbrace{\left(\hat{\theta}_{C_{j}}-\tilde{\theta}_{C_{j}}^{*}(x_{P})\right)}_{T_{1}}+\underbrace{\left(\tilde{\theta}_{C_{j}}^{*}(x_{P})-\mathbb{E}[\tilde{\theta}_{C_{j}}^{*}(x_{P})\mid X\in C_{j}]\right)}_{T_{2}}+\underbrace{\left(\mathbb{E}[\tilde{\theta}_{C_{j}}^{*}(x_{P})\mid X\in C_{j}]-\theta_{C_{j}}^{*}\right)}_{T_{3}}.$
Under standard LLN arguments, the second term satisfies $T_{2}=O_{P}(1/\sqrt{n_{C_{j}}})$, and in an asymptotic regime with $r^{-2}\ll n_{C_{j}}$ one has $T_{2}=o_{P}(r)$. Meanwhile, the first and third terms appear in the proofs of Propositions 2 and 1 of Athey et al. (2019), respectively, and satisfy $T_{1}=o_{P}(r,~{}1/\sqrt{n_{C_{j}}})$ and $T_{3}=\mathcal{O}(r^{2})\implies T_{3}=o(r)$. It follows
$\hat{\theta}_{C_{j}}-\theta_{C_{j}}^{*}=o_{P}\left(r,~{}1/\sqrt{n_{C_{j}}}\right),$
and in particular
$\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}=\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*}+o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right).$
Write $A=\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*}$ and let $E$ be any term satisfying $E=o_{P}(r,~{}1/\sqrt{n_{C_{1}}},~{}1/\sqrt{n_{C_{2}}})$ such that $\Delta(C_{1},C_{2})$ is equivalently written $\Delta(C_{1},C_{2})=(n_{C_{1}}n_{C_{2}}/n_{P}^{2})\cdot\left\lVert A+E\right\rVert^{2}$. Consider the difference
$\Delta(C_{1},C_{2})-\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*}\right\rVert^{2}$ $=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left(\left\lVert A+E\right\rVert^{2}-\left\lVert A\right\rVert^{2}\right),$
$=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left(2\langle A,E\rangle+\left\lVert E\right\rVert^{2}\right).$
Under Specification 2 there exists a fixed proportion $\omega>0$ such that $n_{C_{1}},n_{C_{2}}\geq\omega n_{P}$, and hence $n_{C_{1}}n_{C_{2}}/n_{P}^{2}\geq\omega(1-\omega)$ and also $n_{C_{1}}n_{C_{2}}/n_{P}^{2}\leq 1/4$ for all $n_{C_{1}}+n_{C_{2}}=n_{P}$. Therefore $n_{C_{1}}n_{C_{2}}/n_{P}^{2}=\mathcal{O}(1)$. Meanwhile, $\left\lVert E\right\rVert^{2}=o_{P}(r^{2},~{}1/n_{C_{1}},~{}1/n_{C_{2}})$ is true by definition of $E$, and under our assumptions one may follow the arguments of Athey et al. (2019) Proposition 1 to see that $A=\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*}=\mathcal{O}(r)$. Thus,
$\langle A,E\rangle=\mathcal{O}(r)\cdot o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right)=o_{P}\left(r^{2},~{}\frac{r}{\sqrt{n_{C_{1}}}},~{}\frac{r}{\sqrt{n_{C_{2}}}}\right),$
and therefore
$\Delta(C_{1},C_{2})-\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*}\right\rVert^{2}=o_{P}\left(r^{2},~{}\frac{1}{{n_{C_{1}}}},~{}\frac{1}{{n_{C_{2}}}}\right),$
@@
Suppose the conditions of Lemma B.3 hold, and assume moreover Neyman orthogonal moment conditions such that the underlying Jacobian $V(x)$ defined in Assumption 2 with block form (37). Then,
$\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\eta^{2}\left\lVert V_{\theta\theta}(x_{P})(\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*})\right\rVert^{2}+o_{P}\left(r^{2},~{}\frac{1}{{n_{C_{1}}}},~{}\frac{1}{{n_{C_{2}}}}\right),$
where $\Delta_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ defined in Lemma B.2 denotes the FPT criterion with arbitrary scale factor $\eta\neq 0$.
Proof of Lemma B.4.
From the proof of Lemma B.2 one finds that $\Delta_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ is equivalently written
$\Delta_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})\coloneqq\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\eta^{2}\left\lVert\xi^{\top}(\overline{\psi}_{C_{1}}-\overline{\psi}_{C_{2}})\right\rVert^{2},\qquad\overline{\psi}_{C_{j}}\coloneqq\frac{1}{n_{C_{1}}}\sum_{\{i:X_{i}\in C_{j}\}}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i}).$
Under standard LLN arguments the average scores $\overline{\psi}_{C_{j}}$ satisfy
$\overline{\psi}_{C_{j}}=\mathbb{E}[\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O)\mid X\in C_{j}]+O_{P}(1/\sqrt{n_{C_{j}}}).$ (44)
One applies iterated expectation to see
$\mathbb{E}[\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O)\mid X\in C_{j}]=\mathbb{E}\left[\mathbb{E}\left[\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O)\mid X\right]\mid X\in C_{j}\right]=\mathbb{E}[M_{\hat{\theta}_{P},\hat{\nu}_{P}}(X)\mid X\in C_{j}],$
@@
$\overline{\psi}_{C_{j}}=\mathbb{E}[M_{\hat{\theta}_{P},\hat{\nu}_{P}}(X)\mid X\in C_{j}]+O_{P}(1/\sqrt{n_{C_{j}}}).$ (45)
Expansion of $\bm{M_{\hat{\theta}_{P},\hat{\nu}_{P}}(X)}$.
Under Assumption 2 one considers the Taylor expansion of $M_{\hat{\theta}_{P},\hat{\nu}_{P}}(X)$ about $(\theta,\nu)=(\theta^{*}(x_{P}),\nu^{*}(x_{P}))$:
$M_{\hat{\theta}_{P},\hat{\nu}_{P}}(X)$ $=M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)$
$\hphantom{...}+\left[\nabla_{(\theta,\nu)}M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)\right]\begin{bmatrix}\hat{\theta}_{P}-\theta^{*}(x_{P})\\ \hat{\nu}_{P}-\nu^{*}(x_{P})\end{bmatrix}+O_{P}\left(\left\lVert\begin{bmatrix}\hat{\theta}_{P}-\theta^{*}(x_{P})\\ \hat{\nu}_{P}-\nu^{*}(x_{P})\end{bmatrix}\right\rVert^{2}\right).$
The consistency of the parent solutions $(\hat{\theta}_{P},\hat{\nu}_{P})$ for $(\theta^{*}(x_{P}),\nu^{*}(x_{P}))$ is established by Athey et al. (2019), and in particular $(\hat{\theta}_{P},\hat{\nu}_{P})-(\theta^{*}(x_{P}),\nu^{*}(x_{P}))=O_{P}(r,~{}1/\sqrt{n_{P}})$. The asymptotic regime $r^{-2}\ll n_{P}$ implies $1/\sqrt{n_{P}}=o(r)$ and therefore the higher order quadratic term is equivalently expressed:
$M_{\hat{\theta}_{P},\hat{\nu}_{P}}(X)=M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)+\left[\nabla_{(\theta,\nu)}M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)\right]\begin{bmatrix}\hat{\theta}_{P}-\theta^{*}(x_{P})\\ \hat{\nu}_{P}-\nu^{*}(x_{P})\end{bmatrix}+O_{P}(r^{2}),$
and therefore
$\mathbb{E}\left[M_{\hat{\theta}_{P},\hat{\nu}_{P}}(X)\mid X\in C_{j}\right]$ $=\mathbb{E}\left[M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)\mid X\in C_{j}\right]$
$\hphantom{.....}+\mathbb{E}\left[\nabla_{(\theta,\nu)}M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)\,\middle|\,X\in C_{j}\right]\begin{bmatrix}\hat{\theta}_{P}-\theta^{*}(x_{P})\\ \hat{\nu}_{P}-\nu^{*}(x_{P})\end{bmatrix}+O_{P}(r^{2}).$
One has $\nabla_{(\theta,\nu)}M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)=V(x_{P})+O_{P}(r)$ because $M_{\theta,\nu}(x)$ is Lipschitz in $x$, and the expansion in the previous display becomes:
$\mathbb{E}\left[M_{\hat{\theta}_{P},\hat{\nu}_{P}}(X)\mid X\in C_{j}\right]=\mathbb{E}\left[M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)\mid X\in C_{j}\right]+V(x_{P})\begin{bmatrix}\hat{\theta}_{P}-\theta^{*}(x_{P})\\ \hat{\nu}_{P}-\nu^{*}(x_{P})\end{bmatrix}+O_{P}(r^{2}).$ (46)
Expansion of $\bm{M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)}$.
Following similar arguments, the term $M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)$ is expanded about $(\theta,\nu)=(\theta^{*}(X),\nu^{*}(X))$ as:
$M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)$ $=M_{\theta^{*}(X),\nu^{*}(X)}(X)+V(X)\begin{bmatrix}\theta^{*}(x_{P})-\theta^{*}(X)\\ \nu^{*}(x_{P})-\nu^{*}(X)\end{bmatrix}+O_{P}(r^{2}),$
$=V(X)\begin{bmatrix}\theta^{*}(x_{P})-\theta^{*}(X)\\ \nu^{*}(x_{P})-\nu^{*}(X)\end{bmatrix}+O_{P}(r^{2}),$
where $M_{\theta^{*}(X),\nu^{*}(X)}(X)={\bf 0}$ holds because $(\theta^{*}(X),\nu^{*}(X))$ are defined as satisfying the GRF moment conditions (1) local to $X$. One takes the conditional expectation of the previous display:
$\mathbb{E}\left[M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)\mid X\in C_{j}\right]=\mathbb{E}\left[V(X)\begin{bmatrix}\theta^{*}(x_{P})-\theta^{*}(X)\\ \nu^{*}(x_{P})-\nu^{*}(X)\end{bmatrix}\,\middle|\,X\in C_{j}\right]+O_{P}(r^{2}).$
Whenever $X\in C_{j}$ one has $\left\lVert X-x_{P}\right\rVert=\mathcal{O}(r)$, and the same Lipschitz arguments can be applied to see $V(X)=V(x_{P})+O_{P}(r)$ conditional on $X\in C_{j}$, and the previous display simplifies:
$\mathbb{E}\left[M_{\theta^{*}(x_{P}),\nu^{*}(x_{P})}(X)\mid X\in C_{j}\right]=V(x_{P})\begin{bmatrix}\theta^{*}(x_{P})-\theta_{C_{j}}^{*}\\ \nu^{*}(x_{P})-\nu_{C_{j}}^{*}\end{bmatrix}+O_{P}(r^{2}),$ (47)
where $\theta_{C_{j}}^{*}\coloneqq\mathbb{E}[\theta^{*}(X)\mid X\in C_{j}]$ and $\nu_{C_{j}}^{*}\coloneqq\mathbb{E}[\nu^{*}(X)\mid X\in C_{j}]$. Substitute (47) into the conditional expectation (46):
$\mathbb{E}\left[M_{\hat{\theta}_{P},\hat{\nu}_{P}}(X)\mid X\in C_{j}\right]$ $=V(x_{P})\begin{bmatrix}\theta^{*}(x_{P})-\theta_{C_{j}}^{*}\\ \nu^{*}(x_{P})-\nu_{C_{j}}^{*}\end{bmatrix}+V(x_{P})\begin{bmatrix}\hat{\theta}_{P}-\theta^{*}(x_{P})\\ \hat{\nu}_{P}-\nu^{*}(x_{P})\end{bmatrix}+O_{P}(r^{2}),$
$=V(x_{P})\begin{bmatrix}\hat{\theta}_{P}-\theta_{C_{j}}^{*}\\ \hat{\nu}_{P}-\nu_{C_{j}}^{*}\end{bmatrix}+O_{P}(r^{2}).$
Therefore, the child node score averages $\overline{\psi}_{C_{j}}$ in (45) satisfy
$\overline{\psi}_{C_{j}}=V(x_{P})\begin{bmatrix}\hat{\theta}_{P}-\theta_{C_{j}}^{*}\\ \hat{\nu}_{P}-\nu_{C_{j}}^{*}\end{bmatrix}+O_{P}\left(r^{2},~{}1/\sqrt{n_{C_{j}}}\right),$
and the difference $\overline{\psi}_{C_{1}}-\overline{\psi}_{C_{2}}$ satisfies